# DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition of Human Activities in Smart Homes

### v1

@@ 4.2. Data Pre-Processing
Following previous research into the optimal window size $l$ for classifying activities in CASAS (Krishnan and Cook, 2014), as well Hiremath’s et al.’s prior works in this area (Hiremath et al., 2022), we select $l=20$ for all input windows $W_{i}$. We use sliding windows with a stride of size 1 to get sequence samples that only differ by one reading. Since using all of this input data would yield a overly redundant training dataset (which can lead to over-fitting (Hammerla and Plötz, 2015)), we uniformly sample 10% of these sequences for training. To illustrate how this sampling works, consider the Milan dataset: there are 433,665 distinct sensor readings in total, which results in 433,665 - 19 = 433,646 potential input sequences. After taking a 10% sample, we are left with approximately 43,364 data points. We follow the same methodology at test time on a holdout dataset. When running inference for pattern recognition (see Section 7.1.3), we do not sample the windows and keep all of them to get a continuous representation of each sequence for cluster label change point detection.
Additionally, we adopt two heuristic restrictions on window sequence selection, aiming to more closely mirror potential real-world deployment scenarios. These heuristics enable us to sample a constant number of sensor readings (20 in our case) while still adhering to some common-sense temporal boundaries. Below is a list of heuristics applied during sequence sample generation:
  1. (1)
    Do not create samples across day boundaries: we do not want to form sequences that begin on one day and end on the next one to prevent merging activity patterns that naturally reset with each new day.
  2. (2)
    Do not sample with gaps of more than 30 minutes of inactivity to preserve temporal continuity and ensure that the sampled events reflect a cohesive activity period.
To split the data the training and testing sets, we follow the ”leave one day out” cross-validation strategy: we split the data by random days: 80% of distinct days of data are kept for training, and 20% are set aside for testing. By withholding entire days rather than arbitrary sequences, we reduce the risk of data leakage, where the temporal proximity of training and testing sequences might lead to overly optimistic performance estimates.
@@
Instead of temporal split where we use the first 80% of days for training and the last 20% of days for testing, we chose to select 20% of days at random. This decision was motivated by the fact that, for some datasets, the final portion of the year would coincide with the U.S. holiday season (late December and early January), during which residents’ activities differ substantially from other periods. By randomly sampling the test set days, we mitigate seasonal biases and avoid severely underestimating the pipeline performance.
5. Experimental Evaluation: Clustering Stage
In this section, we evaluate the clustering stage of the DISCOVER pipeline (Steps 0-2) to assess the quality of the clusters generated by our approach. Although the clusters have not yet been assigned DISCOVER labels, examining the clusters helps provide insights into the degree to which they match previous hand-annotated labels from CASAS. To validate cluster performance, we use CASAS as a ground truth oracle to temporarily assign a label to each cluster through majority voting. We then perform a direct comparison with fully supervised techniques, and investigate the impact of varying cluster count, examining the trade-off between granularity, label alignment and annotation costs.
5.1. Setup
We assess the quality of the clusters generated by our approach by comparing them to CASAS labels. To map DISCOVER clusters to CASAS labels, we use majority vote to assign a single label to each cluster based on the CASAS labels of its samples. We then assess these cluster-assigned labels, utilizing CASAS labels as ground truth.
[…]
  • Weighted F1: Emphasizes frequent classes by weighting each class’s F1 score by its support:
[…]
    where $\lvert c\rvert$ is the number of samples belonging to class $c$, and $N$ is the total number of samples.
  • Macro F1: Treats all classes equally by averaging their F1 scores:
[…]
Specifically, we compare against two baselines: DeepCASAS (Demirel et al., 2018) and TDOST (Thukral et al., 2024). DeepCASAS is a fully supervised deep learning algorithm that leverages LSTMs to capture temporal dependencies in the data. In contrast, TDOST converts sensor triggers into natural language descriptions, which provides the model with rich contextual information and enables the transfer of a single model across multiple households.
Note that both DeepCASAS and TDOST, as originally published, were trained and evaluated on pre-segmented datasets. This allowed the authors to compare the isolated algorithmic performance in a controlled setting, but is a less realistic scenario compared to operating over continuous sensor streams. To achieve a fair comparison to our approach, which does not assume pre-segmentation, we re-train both DeepCASAS and TDOST algorithms on sliding windows $W_{i}$ of length $l$ to match our sequences. Additionally, given the prevalence of the Other label in our datasets (see Section 4.1), we conduct experiments both with and without the Other label. Specifically, we re-train DeepCASAS and TDOST with and without the Other label, which allows us to evaluate how well these methods work with only relevant activities.
5.2. Results
F1 Weighted
Milan Aruba Cairo Model All Labels Excl. Other All Labels Excl. Other All Labels Excl. Other DeepCASAS 0.61 0.79 0.78 0.93 0.72 0.82 TDOST 0.62 0.83 0.80 0.92 0.74 0.82 DISCOVER 0.53 0.82 0.75 0.90 0.64 0.85
F1 Macro
Milan Aruba Cairo Model All Labels Excl. Other All Labels Excl. Other All Labels Excl. Other DeepCASAS 0.38 0.53 0.48 0.69 0.16 0.46 TDOST 0.40 0.63 0.54 0.69 0.17 0.41 DISCOVER 0.32 0.50 0.24 0.49 0.12 0.39
Table 2. Weighted and macro F1 scores for DeepCASAS, TDOST, and DISCOVER trained and tested on the Milan, Aruba, and Cairo datasets using CASAS labels as ground truth. The best result for each configuration is highlighted in green. While fully supervised methods yield the highest overall performance, the gap in DISCOVER performance is moderate—with comparable F1 scores achieved when data associated with the Other label is excluded, signifying a reasonable alignment between DISCOVER clusters and CASAS labels.
In this section we present how well our clusters align with the CASAS labels, and also experiment with the optimal number of clusters needed to achieve useful performance.
5.2.1. DISCOVER Cluster Alignment with CASAS Labels
[…]
5.2.2. Comparison to Supervised Baselines
[…]
5.2.3. tSNE Analysis
Figure 4. tSNE projection of SCAN embeddings from the Milan household; each point represents a sensor window embedding colored by its original CASAS label. Insets display the deployment environment layout overlaid with a heatmap of sensor activations. Insets (a) and (b) highlight two distinct clusters within the CASAS Cook label—cluster 16 showing movement between kitchen and dining areas, and cluster 5 capturing activity near the medicine cabinet. Insets (c) and (d) show clusters from the Relax label, corresponding to sitting in the TV room armchair and sitting in the living room armchair. Data associated with the Other label (gray) is dispersed across clusters, underscoring its heterogeneous nature. This figure showcases DISCOVER’s capability to uncover more granular and nuanced activity categories than the original CASAS labels.
In this section, we perform a visual inspection of Milan clusters in order to gain more insights into the generated clusters. Fig. 4 shows a tSNE projection of the SCAN embeddings from the Milan dataset; each point represents a 768-dimensional sequence embedding colored by its CASAS label. While some activities (e.g., Work) form single clusters, data representing other activity labels (e.g., Cook, Relax) fall into multiple clusters. Each point in the scatterplot is annotated with a probability score reflecting SCAN’s confidence in assigning that sequence to its respective cluster, offering interpretability and a level of control that simpler clustering methods (e.g., k-means) lack.
Further, we take the two largest Milan activity labels, Cook and Relax, and inspect the data within individual clusters, as shown in the figure insets. Each inset includes the layout of the deployment environment overlaid with a heatmap representing sensor activation patterns for the data sequences associated with that cluster. Fig. 4 illustrates that clustering using DISCOVER facilitates the discovery and differentiation of granular sub-activities. Insets (a) and (b) are associated with the Cook CASAS label (orange). Although both sets of data are labeled as Cook within CASAS, we observe clear differences between the sensor patterns within each cluster. Cluster 16 (inset (a)) captures the person leaving the kitchen and moving to the dining room, with motion sensors from the dining room and part of the dining room registering the person’s presence. By comparison, cluster 5 (inset (b)) is tightly focused on activity around the medicine cabinet in the kitchen. Insets (c) and (d) are associated with the Relax CASAS label (burgundy), where cluster 3 (inset (c)) captures a static activity in the TV room armchair while cluster 17 (inset (d)) corresponds to sitting in a living room armchair instead.
Note that data associated with the Other label (gray) is scattered throughout multiple clusters in Fig. 4, indicating that it does not represent a single sub-activity. Instead, Other activities appear to correspond to multiple sub-activities.
[…]
5.2.4. Impact of the Number of Clusters
Figure 5. Macro F1 scores for varying numbers of SCAN clusters ($k$) on Milan, Aruba, and Cairo datasets, using all CASAS labels in (a), and without the Other label in (b). We chart shows average F1 scores with bootstrapped 95% confidence intervals . Increasing $k$ up until 20-40 clusters improves alignment with CASAS labels before the performance improvement stagnates, suggesting that the optimal number of clusters lies in that range.
We also investigate how the choice of $k$—the number of clusters—impacts our model’s ability to capture and distinguish fine-grained activities. Similar to previous section, we use CASAS labels and apply majority voting per cluster to compute Macro F1 scores.
Fig. 5 shows the Macro F1 scores with bootstrapped 95% confidence intervals (CIs) for Milan, Aruba, and Cairo, when $k\in\{10,15,20,30,40,50,60,100\}$ clusters. The left-hand panels in (a) show performance when all CASAS labels are considered; the right-hand panels in (b) exclude Other. These F1 scores indicate that utilizing more clusters generally leads to better classification performance in (a) for all three datasets, although the improvements appear to plateau after $k=30$ for Milan, and $k=20$ for Aruba and Cairo. When we exclude the Other label, the macro F1 scores for both Milan and Aruba plateau when $k>30$ and further increases are not statistically significant; Cairo stays relatively flat after $k=20$, and the improvements for $k>=50$ have very wide CIs. This leads to a conclusion that increasing the number of clusters beyond 20–30 is not particularly useful. Refer to Fig. 10 in the Appendix to see what these various clusters look like on 2D tSNE projections.
, as the F1 scores plateau. In addition, we also visualizing the effect of the number of clusters through a tSNE plot in Fig. 10 in the Appendix.
6. Experimental Evaluation: Labeling Stage
For our second set of experiments, we move beyond the CASAS annotations and evaluate the labeling stage of our pipeline (step 2, 3, and 4 in 1). This resembles the envisioned use case for our approach where we employ our active learning-like scheme, driven by our self-supervised clustering stage, to guide human annotator to provide reliable labels thereby minimizing manual efforts.
By comparing our human-labeled clusters to CASAS labels, we highlight how our approach provides a more granular and detailed categorization of (sub-)activities, offering finer resolution and capturing nuances that the original labels may miss. We also evaluate the quality and consistency of these labels by examining two key metrics: inter-rater agreement, which measures consistency between different labelers, and cluster agreement, which assesses the precision of the labels assigned to each cluster.
6.1. Setup
Our objective is to present human labelers with visualizations of a small number of representative data sequences from each cluster, and for labelers to select the appropriate label for each sample. Below, we detail the labels selected for our work and evaluation metrics.
6.1.1. Label Set
The DISCOVER labeling scheme is organized hierarchically to accommodate different levels of detail. At the highest level, we organize all activities into categories such as single-room and multi-room events, with subcategories specific to each environment (e.g., kitchen vs. bedroom). By allowing annotators to choose increasingly granular labels, we preserve the flexibility needed for diverse research objectives.
@@ 6.1.1. Label Set
To capture a finer level of granularity than conventional HAR labels such as Cook, Relax, or Work, we developed a set of sub-activity labels by studying the physical layouts of the CASAS homes, examining the replays of actual sensor activations throughout the day, and drawing insights from prior work on structural constructs in activities (Hiremath and Plötz, 2024). We selected a large number of potential labels in order to avoid over-constraining our annotators, leveraging information about furniture placement when it was available. Our labels capture both activities within a single room (e.g., movement all over kitchen, movement near fridge, movement near stove, and movement near medicine cabinet) and movement actions between the rooms (e.g., walking from bedroom to office, leaving guest bathroom). See the full list of sub-activities and their hierarchies in Fig. 11 in the Appendix.
Note, we have chosen a dataset-specific approach in our label selection, with several labels tailored specifically to the CASAS dataset (e.g., medicine cabinet), in order to demonstrate the degree to which DISCOVER supports customization to a given environment or use case. The granularity of the labels is fully adjustable; depending on the specific application, a user can, for example, choose whether they care to split armchair sitting and couch sitting into distinct classes, or if they want to group them into a broader Relax category.
6.1.2. Labeling Sample Selection and Metrics
@@
To obtain cluster labels, we recruited 15 independent raters to each individually label data segments. For each cluster $c_{k}$, we choose $m$ sequence samples that are closest to the cluster centroid for labeling, with $m=5$ in this work. With $k=20$ clusters per model, this resulted in a total of $300$ samples across all three datasets. Samples were randomized and each sample was labeled by two raters, resulting in $600$ total sample-label pairs. To generate a label, each rater was presented with a selection of $40$ samples, for which they assigned labels using the web-based DISCOVER annotation tool. Fig. 3 shows a screenshot of a sample annotation replay for the Milan household. During the replay, the rater can view a temporal playback of the sensor activations captured by the data sequence, and use a drop-down menu to select a label. Note that the DISCOVER tool is fully customizable and is available in open source at https://anonymized.
@@
Since our sub-activity labels follow a hierarchical, multi-level taxonomy (see Fig. 11), disagreements may sometimes stem from different levels of specificity (e.g., movement near fridge vs. movement in kitchen). To account for this, we also compute Kappa scores at a coarser level by “leveling up” each annotated label to its parent category. For instance, if one rater selects movement near fridge and the other chooses movement in kitchen, these would be treated as the same label at the higher level. Comparing Kappa scores at different levels of the hierarchy helps distinguish minor discrepancies in labeling granularity from dramatically different activity interpretations (e.g., bathroom activity vs. reading in armchair).
6.2. Results
@@
Table 3. Sub-Activity label quality: inter-rater (Cohen’s) and cluster-agreement (Fleiss’s) Kappa scores for each of the three CASAS datasets (Milan, Aruba, and Cairo). “Original Label” indicates the use of the most specific labels chosen by each rater (e.g., movement near fridge), while “Level-Up Label” merges them into coarser categories (e.g., movement in kitchen). Overall, very high inter-rater agreement indicates very high alignment between annotators; decent cluster agreement also implies relative homogeneity among the ratings. The resulting increase in Kappa scores from “Original Label” to “Level-Up Label” highlights that most disagreements stem from label specificity rather than fundamentally different interpretations of the underlying activity.
In this section, we first analyze the inter-rater and cluster-agreement scores, then discuss how the obtained labels different from CASAS labels. Table 3 summarizes the inter-rater and cluster-agreement scores for Milan, Aruba, and Cairo. We also refer to Fig. 12 in the Appendix for a detailed table showing the majority label assigned to each cluster and its corresponding vote count.
Inter-rater agreement, measured as Cohen’s Kappa, exceeds 0.85 for all three datasets (Milan $\kappa=0.85$, Aruba $\kappa=0.89$, Cairo $\kappa=0.87$). These values indicate very high consistency between the two independent annotators in how they interpreted and labeled the same replay samples. Furthermore, when we “level up” each label to its parent category in our hierarchical taxonomy (e.g., merging movement near bed into movement in bedroom), the Kappa scores rise even further—reaching 0.88 in Milan and over 0.90 in Aruba and Cairo. This gain reflects the fact that any disagreements that did arise between raters, were largely in relation to differences in label specificity, such as sitting in armchair vs. movement in living room, rather than fundamentally diverging perceptions of the activity itself.
Cluster agreement, measured as Fleiss’ Kappa, results in moderately high scores ranging from 0.60–0.63 when annotators used highly specific sub-activity labels. Once labels are rolled up to a coarser levels, these scores climb to 0.75–0.81 across the three datasets. Here again, the improvement underscores that most labeling discrepancies within a cluster stem from variation in granularity rather than actual activity disagreements. For instance, in one cluster designated kitchen activity, half of the annotators specified movement near medicine cabinet, while the other half used movement in kitchen. In another cluster, some annotators perceived simultaneous sensor firings in two rooms as multi-room activity, while others focused on whichever room had the most events.

### v3

@@ 4.2. Data Pre-Processing
Following previous research into the optimal window size $l$ for classifying activities in CASAS (Krishnan and Cook, 2014), we select $l=20$ for all input windows $W_{i}$. We use sliding windows with a stride of size 1 to get sequence samples that only differ by one reading. Since using all of this input data would yield an overly redundant training dataset (which can lead to over-fitting (Hammerla and Plötz, 2015)), we uniformly sample 10% of these sequences for training. For example, consider the Milan dataset: there are 433,665 distinct sensor readings in total, which results in 433,665 - 19 = 433,646 potential input sequences. After taking a 10% sample, we are left with approximately 43,364 data points.
@@
To split the data into training and testing sets, we follow the “leave one day out” cross-validation strategy: we split the data by random days: 80% of distinct days of data are kept for training, and 20% are set aside for testing. By withholding entire days rather than arbitrary sequences, we reduce the risk of data leakage, where the temporal proximity of training and testing sequences might lead to overly optimistic performance estimates. Instead of temporal split where we use the first 80% of days for training and the last 20% of days for testing, we chose to select 20% of days at random. This decision was motivated by the fact that, for some datasets, the final portion of the year would coincide with the U.S. holiday season (late December and early January), during which residents’ activities differ substantially from other periods. By randomly sampling the test set days, we mitigate seasonal biases and avoid severely underestimating the pipeline performance.
4.3. Defining a Flexible Taxonomy for Pattern Interpretation
Our final design choice concerns the selection of labels used to interpret the clusters discovered by DISCOVER, which are used in the drop-down menu in Figure 3. While the inclusion of a label set may appear counterintuitive in a pattern-centric framework, it is important to emphasize that DISCOVER labels are fundamentally different from predefined activity labels commonly used in supervised HAR. Rather than representing a fixed, closed vocabulary of activities that the data must conform to, our labels define a broad space of possible behavioral patterns, many of which may not occur in a given dataset. Moreover, sensor data are never pre-labeled with these categories; instead, labels are assigned only after unsupervised pattern discovery, making the approach weakly supervised rather than fully supervised. In practice, DISCOVER does not require a closed label set at all, labels are introduced here primarily to support systematic analysis and comparison.
@@ 4.3. Defining a Flexible Taxonomy for Pattern Interpretation
Within this paper, we organize DISCOVER labels hierarchically to support flexible interpretation at multiple levels of abstraction. At the highest level, patterns are grouped into broad categories of single-room and multi-room events, with progressively finer-grained subcategories tailored to each environment (e.g., kitchen versus bedroom). This hierarchical structure allows annotators to select labels at the level of granularity most appropriate for the observed behavior pattern. We include both conventional HAR labels (e.g., Cook, Relax, Work) and as well as provided more detailed options by studying the physical layouts of the CASAS homes and drawing insights from prior work on structural constructs in activities (Hiremath and Plötz, 2024). Examples include activities within a single room (e.g., movement all over kitchen, movement near fridge, movement near stove, and movement near medicine cabinet) and movement actions between the rooms (e.g., walking from bedroom to office, leaving guest bathroom). See the full list of patterns of daily living and their hierarchies in Fig. 11 in the Appendix.
5. Results: Reliability and Coherence of Discovered PDLs
In this section, we evaluate the patterns of daily living identified by DISCOVER and compare our human-labeled clusters to CASAS labels to highlight how our approach provides a more fine-grained and detailed categorization of activities, offering finer resolution and capturing nuances that the original labels may miss. We also evaluate the quality and consistency of the DISCOVER labels by examining two key metrics: inter-rater agreement, which measures consistency between different labelers, and cluster agreement, which assesses the precision of the labels assigned to each cluster.
5.1. The Annotation Protocol
@@
To identify patterns of daily living, we generated $k=20$ clusters per dataset using the steps outlined in Sections 3.1 and 3.2. To obtain a label for each cluster $c_{k}$, we then chose $m$ sequence samples that are closest to the cluster centroid for labeling, with $m=5$ in this work. This resulted in a total of $300$ samples across all three datasets. Samples were randomized and each sample was labeled by two raters, resulting in $600$ total sample-label pairs. We recruited 15 raters, including five graduate students and ten working professionals. Participants had no prior experience with ambient sensor data; thus, their labeling decisions were based solely on the visual representations provided by the DISCOVER annotation tool (Karpekov et al., 2025). Fig. 3 shows a screenshot of a sample annotation replay for the Milan household. During the replay, the rater can view a temporal playback of the sensor activations captured by the data sequence, and use a drop-down menu to select a label, or choose “Other” if none of the labels apply. Note that the DISCOVER tool is fully customizable and is available in open source at https://anonymized.
5.2. Evaluation of Consensus and Cluster Homogeneity
@@
Since our labels follow a hierarchical, multi-level taxonomy (see Fig. 11 in the Appendix), disagreements may sometimes stem from different levels of specificity (e.g., movement near fridge vs. movement in kitchen). To account for this, we also compute Kappa scores at a coarser level by “leveling up” each annotated label to its parent category. For instance, if one rater selects movement near fridge and the other chooses movement in kitchen, these would be treated as the same label at the higher level. Comparing Kappa scores at different levels of the hierarchy helps distinguish minor discrepancies in labeling granularity from dramatically different activity interpretations (e.g., bathroom activity vs. reading in armchair).
@@
Table 2. Patterns of Daily Living label quality: inter-rater (Cohen’s) and cluster-agreement (Fleiss’s) Kappa scores for each of the three CASAS datasets (Milan, Aruba, and Cairo). “Original Label” indicates the use of the most specific labels chosen by each rater (e.g., movement near fridge), while “Level-Up Label” merges them into coarser categories (e.g., movement in kitchen). Overall, very high inter-rater agreement indicates very high alignment between annotators; decent cluster agreement also implies relative homogeneity among the ratings. The resulting increase in Kappa scores from “Original Label” to “Level-Up Label” highlights that most disagreements stem from label specificity rather than fundamentally different interpretations of the underlying activity.
In this section, we first analyze the inter-rater and cluster-agreement scores, then discuss how the obtained labels differ from CASAS labels. Table 2 summarizes the inter-rater and cluster-agreement scores for Milan, Aruba, and Cairo. We also refer to Fig. 12 in the Appendix for a detailed table showing the majority label assigned to each cluster and its corresponding vote count.
5.2.1. Inter-Rater Reliability.
Inter-rater agreement, measured as Cohen’s Kappa, exceeds 0.85 for all three datasets: Milan $\kappa=0.85$, Aruba $\kappa=0.89$, Cairo $\kappa=0.87$. These values indicate very high consistency between the two independent annotators in how they interpreted and labeled the same replay samples. Furthermore, when we “level up” each label to its parent category in our hierarchical taxonomy (e.g., merging movement near bed into movement in bedroom), the Kappa scores rise even further—reaching 0.88 in Milan and over 0.90 in Aruba and Cairo. This gain reflects the fact that any disagreements that did arise between raters were largely in relation to differences in label specificity, such as sitting in armchair vs. movement in living room, rather than fundamentally diverging perceptions of the activity itself.
5.2.2. Cluster Homogeneity.
Cluster agreement, measured as Fleiss’ Kappa, results in moderately high scores ranging from 0.60–0.63 when annotators used highly specific patterns of daily living labels. Once labels are rolled up to a coarser levels, these scores climb to 0.75–0.81 across the three datasets. High values of Fleiss’ Kappa highlights cluster homogeneity, demonstrating that discovered clusters correspond to coherent and semantically consistent patterns of daily living rather than arbitrary groupings of sensor windows. The improvement underscores that most labeling discrepancies within a cluster stem from variation in granularity rather than actual activity disagreements. For instance, in one cluster designated kitchen activity, half of the annotators specified movement near medicine cabinet, while the other half used movement in kitchen. In another cluster, some annotators perceived simultaneous sensor firings in two rooms as multi-room activity, while others focused on whichever room had the most events.